OpenStudy (camijacobson):

What is a solution to the equation y=3x+1? A. (3,10) B. (2,6) C. (1,5) D. (0,3)

OpenStudy (anonymous):

i think its D

jimthompson5910 (jim_thompson5910):

Let's plug in a random answer choice. I'm going to pick choice C (1,5) means x = 1 and y = 5 y=3x+1 5=3*1+1 ... replace x with 1, replace y with 5 5 = 3+1 5 = 4 ... this is false Since "5 = 4" is false, this means y=3x+1 is false when (x,y) = (1,5) So (1,5) is NOT a solution. If it were a solution, then it would make y=3x+1 true All of this allows us to rule out choice C. Do the same for the other choices.

OpenStudy (anonymous):

A